OpenStudy (anonymous):

find the equation for the line that passes through the point (1,-3) and that is parallel to the line with the equation (y-2)=1/4 (x+4)

OpenStudy (anonymous):

parallel means same slope, so the slop is 1/4

OpenStudy (anonymous):

next just plug the point (1,-3) into a point-slope equation format and ur done!

OpenStudy (anonymous):

How would I be able to do that?

OpenStudy (anonymous):

(y-2)=1/4 (x+4) its the same format as this

OpenStudy (anonymous):

(y-y1)=m (x+x1) where m = the slop which is 1/4. (x1, y1) is the coordinate of ur point (1,-3)

OpenStudy (anonymous):

Is the answer y=1/4x+13/4?

OpenStudy (anonymous):

u mean y=1/4x - 13/4?

OpenStudy (anonymous):

y-(-3) = y+3

OpenStudy (anonymous):

so you would subtract 3 from both sides

OpenStudy (anonymous):

Okay that makes sense. Thank You.

OpenStudy (anonymous):

no problem